# A High-Quality Dataset and Reliable Evaluation for Interleaved Image-Text Generation

### v1

@@ 3.4 Dataset Composition
InterSyn contains approximately 1.8 million single-turn samples and 50k multi-turn dialogues. Data quality is ensured through the SEIR method, which iteratively refines samples to improve answer accuracy, conversational coherence, and image-text synergy. InterSyn offers a rare combination of diversity and quality, providing a robust foundation for training unified multimodal models with strong instruction-following and contextual reasoning capabilities.
@@ 4 Experiment
4.1 Evaluated Models and Evaluation Methods
4.1.1 Evaluated Models
@@
We evaluate 11 multimodal generators $G$ capable of producing interleaved image-text outputs, covering a representative spectrum of existing models. For comparative analysis, we categorize them into two groups based on whether they natively support interleaved generation within a single stage:
@@
Non-Interleaved Generators: These generators produce text and images sequentially through modular pipelines, which include: Emu3 [31], Janus-Pro [7], VILA-U [30], Show-o [29], Show-o-Turbo [43], Liquid [44], D-DiT [45], GPT-4o [46] + DALL·E3 [5], and Gemini1.5 [47] + FLUX [48].
@@
Interleaved Generators: These generators are designed to generate interleaved text-image outputs within a unified process, including VARGPT [49] and Anole [32].
4.1.2 Evaluation Methods
For evaluation, we adopt the benchmark introduced in Section 3.2. These questions are used to prompt all generators. In addition, to validate the effectiveness of the SEIR method, we also generate responses using different refinement iterations. Each generated response is independently evaluated by the following judges across four dimensions: TCC, ICC, IQ and ITS.
Human Judge. Human evaluation is conducted first. To ensure scoring reliability, we adopt a cross-review protocol: when significant discrepancies arise between annotators, discussions are held to mitigate individual bias and enhance evaluation quality.
Model-based Judge. We also use multimodal large language models (MLLMs) for automated assessment. Specifically, GPT-4o [46], InternVL2.5 [2], and QwenVL2.5 [25] are employed to score the generated outputs. While model-based evaluation significantly improves efficiency, it may deviate from human judgment on tasks requiring complex reasoning or fine-grained understanding.
SynJudge.
To ensure both accuracy and efficiency, we introduce SynJudge, an automated scoring model optimized for human consistency and scalability. SynJudge is trained on 38,400 human annotated samples, with QwenVL2.5 [25] and InternVL2.5 [2] explored as backbone models.
Finally, we obtain five model-based judges $M$ and a human judge $H$ consisting of ten experts.
4.2 Evaluation Metrics
To quantitatively assess the performance and consistency of judges and generators, we adopt the following metrics:
Mean Score. Measures the average quality of a generator’s outputs under each evaluation dimension.
Score Variance. Captures the consistency of a generator’s outputs across different questions; lower variance indicates more stable performance.
Root Mean Squared Error (RMSE) Quantifies the deviation between a model-based judge’s scores and human scores; lower RMSE implies better alignment with human judgment.
Formal definitions of these metrics are provided in the Appendix.
4.3 Experimental Results
4.3.1 Question Quality
We conduct human evaluation to assess the impact of QR. As shown in Figure 3, quality improves over the first three iterations but plateaus thereafter. Based on this, we fix the number of refinements to 3, achieving 99.5% of peak quality while reducing computational cost by 40%.
@@ 4.3.1 Question Quality
Table 1: Impact of answer refinement (AR) and image refinement (IR) on answer quality. The table reports human evaluation scores across four dimensions (TCC, ICC, IQ, ITS). AR improves TCC, ICC, and ITS, while IR further enhances ICC and ITS, confirming the effectiveness of iterative refinement.
@@
4.3.2 Answer Quality
We conduct human evaluation across different iterations of AR and IR, using a set of 500 prototype questions generated through three rounds of QR. As shown in Table 4.3.1, the results confirm that AR primarily improves content informativeness, while IR enhances multimodal synergy, demonstrating the effectiveness of the SEIR method. Experimental results show that when both AR and IR are set to 4 or 5, the improvements become marginal. Based on this, we fix the number of AR and IR iterations to 3 in dataset construction settings.
4.3.3 Judges Comparison
Figure 4: Visualization of RMSE across evaluation dimensions for different judges.
To verify the scoring ability of different judges, we compute the RMSE between five model-based judges and human across four evaluation dimensions, using a human-annotated test set of 9,600 items. Table 2 and Figure 4 present the numerical results and their corresponding visualizations. The radar plot shows that the smaller its range, the better the performance of the judge. Among all judges, QwenVL-trained shows the closest alignment with human judgments and is therefore selected as the final judge, referred to as SynJudge.
Table 2: RMSE Comparison of different judges across four evaluation dimensions
Generator GPT-4o QwenVL InternVL QwenVL_trained InternVL_trained
TCC ICC IQ ITS TCC ICC IQ ITS TCC ICC IQ ITS TCC ICC IQ ITS TCC ICC IQ ITS
Anole 1.16 1.10 1.13 1.18 0.56 1.34 0.99 1.37 1.12 1.04 1.22 1.04 0.47 0.75 0.81 0.75 0.45 0.79 0.86 0.72
DDiT 0.61 1.12 1.11 1.34 0.59 1.10 1.12 0.66 0.59 0.91 1.10 0.67 0.33 0.67 0.68 0.35 0.41 0.68 0.77 0.44
Emu3 0.96 1.11 0.90 1.24 0.76 1.32 0.81 1.22 0.86 0.95 1.00 1.11 0.58 0.70 0.61 0.68 0.57 0.75 0.77 0.84
InterSyn 0.88 0.84 0.74 1.15 0.87 0.89 0.80 1.23 0.82 0.75 0.75 1.04 0.63 0.71 0.60 0.65 0.63 0.56 0.76 0.66
Gemini+Flux 0.82 0.79 0.68 0.90 0.78 1.16 0.76 0.89 0.76 0.73 0.70 0.91 0.64 0.67 0.61 0.67 0.60 0.65 0.57 0.73
GPT-4o+DALL-E3 0.67 0.73 0.80 0.85 0.62 0.74 0.90 1.08 0.63 0.72 0.84 0.87 0.50 0.59 0.55 0.65 0.47 0.58 0.58 0.61
Janus-Pro 1.31 1.26 1.19 1.38 1.18 1.46 1.27 1.37 1.32 1.05 1.41 1.23 0.47 0.77 0.81 0.79 0.43 0.73 0.83 0.90
Liquid 1.06 1.07 1.05 1.40 0.86 1.00 0.81 1.74 1.14 0.98 1.20 1.33 0.49 0.77 0.76 0.78 0.48 0.64 0.74 0.82
Show-o 1.02 1.19 1.14 1.33 0.96 1.10 1.16 1.23 0.89 1.00 1.20 1.24 0.61 0.79 0.67 0.77 0.65 0.77 0.75 0.94
Show-o-Turbo 1.25 1.09 1.14 1.25 1.22 1.09 0.94 1.43 1.00 0.94 1.21 1.16 0.77 0.81 0.82 0.71 0.77 0.83 0.76 0.75
VARGPT 1.37 0.81 0.62 0.77 0.71 0.80 0.78 1.00 1.52 0.74 0.71 0.75 0.52 0.51 0.41 0.56 0.63 0.64 0.61 0.61
VILA-U 1.00 1.17 1.28 1.39 0.59 1.12 1.20 1.13 0.85 1.02 1.40 0.98 0.42 0.87 0.81 0.71 0.45 0.76 0.82 0.62
4.3.4 Generator evaluation
@@
We evaluate the responses of 11 generators alongside those produced by our SEIR pipeline using the benchmark introduced in Sec 3.2. Both human judge and SynJudge agree that SEIR-generated samples (InterSyn) achieve the highest mean scores across all evaluation dimensions, outperforming the strongest baseline, GPT-4o+DALL-E, by 0.34–0.66; the largest margin appears in ITS. InterSyn also shows the low variances (below 0.61), reflecting consistently high-quality outputs and underscoring the robustness of our automatically constructed dataset. Crucially, the sizeable margin that remains between InterSyn and the best existing generator reveals that current state-of-the-art models still struggle with image–text alignment and complementarity, leaving substantial room for future improvement.
4.3.5 Generative Model Fine-Tuning
Table 3: Score Comparison after Fine-Tuning.
Model TCC ICC IQ ITS
Anole 3.09 3.01 2.92 2.26
Anole_trained 3.52 3.24 3.1 2.94
VILA-U 2.46 3.72 3.37 2.19
VILA-U_trained 3.19 3.83 3.39 3.33
To verify the effectiveness of the InterSyn dataset, we fine-tune Anole and VILA-U on 50K samples from InterSyn. As shown in Table 3, fine-tuning leads to overall performance improvements across all evaluation dimensions. Specifically, gains in IQ remain limited, likely due to intrinsic constraints in the models’ visual generation capabilities. In contrast, more substantial improvements are observed in the other three metrics—particularly TCC and ITS—highlighting the InterSyn dataset’s effectiveness in enhancing both semantic alignment and answer completeness.
5 Ablation Study

### v2

@@ 3.4 InterSyn Composition
InterSyn contains approximately 1.8 million single-turn samples and 50,000 multi-turn dialogues. Data quality is ensured through the SEIR method, which iteratively refines samples to improve answer accuracy, conversational coherence, and image-text synergy. InterSyn offers a rare combination of diversity and quality, providing a robust foundation for training unified multimodal models with strong instruction-following and contextual reasoning capabilities.
3.5 SynJudge: A Reliable Evaluator for Interleaved Outputs
To reliably evaluate the complex, instruction-following capabilities of interleaved image-text generators, we propose SynJudge, a reliable and comprehensive judge model that demonstrates high alignment with human judgment.
3.5.1 Evaluation Dimensions
SynJudge is designed to provide interpretable, quantitative feedback by assessing responses across four key, complementary dimensions (See the appendix §F for details.):
Text Content Completeness (TCC): Assesses whether the generated text completely and accurately answers the user’s question, covering all sub-tasks and constraints.
Image Content Completeness (ICC): Assesses whether the generated images correctly depict the key objects, scenes, or concepts requested in the prompt.
Image Quality (IQ): Evaluates the visual fidelity, aesthetics, and overall quality of the generated images, penalizing artifacts, distortions, or low resolution.
Image–Text Synergy (ITS): Measures the cross-modal relationship. Unlike simple consistency metrics, ITS specifically rewards tight, complementary alignment where the text and images work together to form a cohesive answer. It penalizes redundancy (e.g., text merely describing the image) and irrelevance.
3.5.2 Training and Model Selection
To ensure robust evaluation, we first constructed a high-quality, human-annotated dataset. The process began by collecting a diverse set of interleaved responses from various generators (detailed in §4.1.2). These responses were then rigorously scored by a panel of ten expert annotators across the four dimensions (TCC, ICC, IQ, ITS). To ensure high reliability, each sample was scored by multiple experts. Any scoring inconsistencies were resolved through a mandatory discussion protocol until a unified, final score was reached for each sample. This effort yielded a total of 48,000 annotated pairs, which were split into a 38,400-sample training set and a held-out 9,600-sample validation set. Using this data, we fine-tuned strong MLLM backbones, including QwenVL and InternVL. As demonstrated in our experiments (see §4.4), the model trained from QwenVL achieved the strongest alignment with human judgment, exhibiting the lowest RMSE and highest agreement (A@1). Consequently, we designate this QwenVL-trained model as our final SynJudge.
3.5.3 Evaluation Benchmark for Generators
To rigorously assess the quality of interleaved image-text outputs, we established a fixed Evaluation Benchmark. This benchmark comprises a curated series of 4,000 questions (500 human-authored from §3.2 and 3,500 SEIR-generated) spanning the full topic hierarchy. By requiring different models to generate responses to this identical set of questions, we ensure a fair, standardized, and comprehensive comparison of their instruction-following and multimodal generation capabilities.
@@ 4 Experiment
4.1 Experimental Setup
4.1.1 Evaluation Dimensions for Interleaved Outputs
We adopt the four evaluation dimensions defined in §3.5.1: TCC, ICC, IQ, and ITS. Briefly, TCC and ICC assess content correctness in text and image respectively; IQ focuses on visual fidelity; and ITS specifically measures the cross-modal synergy and alignment.
To quantitatively assess generators’ performance, we adopt the mean score (average score across every dimension) and variance score (stability across questions). See the appendix §D.2 for details.
4.1.2 Evaluated Generators
@@
We evaluate 13 multimodal generators $G$ capable of producing interleaved image-text outputs and we categorize them into two groups based on whether they natively support interleaved generation:
@@
(1) Non-Interleaved Generators: These generators produce text and images sequentially through modular pipelines, which include: Emu3 (Wang et al., 2024b), Janus-Pro (Chen et al., 2025b), VILA-U (Wu et al., 2024), Show-o (Xie et al., 2024), Show-o-Turbo (Xu et al., 2025), Liquid (Wu et al., 2025), D-DiT (Li et al., 2025), GPT-4o (OpenAI, 2024) + DALL-E3 (Betker et al., 2023), and Gemini2.5 (Comanici et al., 2025) + FLUX (Black Forest Labs, 2024).
@@
(2) Interleaved Generators: These generators can generate interleaved image-text outputs within a unified process, including VARGPT (Zhuang et al., 2025a), VARGPT-v1.1 (Zhuang et al., 2025b), Anole (Chern et al., 2024) and BAGEL (Deng et al., 2025).
4.2 Efficiency of InterSyn
4.2.1 Data Efficiency and Scalability
We fine-tune four generators on randomly sampled InterSyn subsets of sizes 25k/50k/100k/200k and evaluate with SynJudge. Table 1 shows consistent improvements as data increases. Notably, even 25k/50k samples already yield clear gains across all dimensions, and further scaling to 200k continues to improve TCC, ICC, and especially ITS, which highlights the InterSyn dataset’s effectiveness in enhancing both semantic alignment and answer completeness.
Table 1: Fine-tuning results on varying subset sizes of InterSyn. Performance consistently improves as training data scales from 25K to 200K samples, demonstrating the dataset’s effectiveness and scalability. Notably, just 50K samples yield substantial gains across all models, with continued improvement in content and synergy metrics (TCC, ICC, ITS) at larger scales. All scores are SynJudge means.
Anole VILA-U VARGPT-v1.1 BAGEL
Subset TCC ICC IQ ITS TCC ICC IQ ITS TCC ICC IQ ITS TCC ICC IQ ITS
baseline 3.09 3.01 2.92 2.26 2.46 3.72 3.37 2.19 3.26 1.01 1.23 0.68 3.11 3.89 4.23 2.87
+ 25k 3.35 3.25 3.01 2.40 2.95 3.78 3.38 2.90 3.51 2.45 2.90 2.55 3.45 4.02 4.21 3.28
+ 50k 3.47 3.28 3.10 2.74 3.19 3.83 3.39 3.20 3.68 3.12 3.67 3.00 3.69 4.11 4.19 3.56
+ 100k 3.51 3.41 3.13 2.87 3.33 3.88 3.31 3.28 3.73 3.22 3.66 3.20 3.87 4.09 4.31 3.78
+ 200k 3.64 3.52 3.08 3.11 3.49 4.01 3.40 3.47 3.86 3.39 3.72 3.53 4.13 4.18 4.25 4.02
4.2.2 Retention of General Understanding
Table 2: Understanding performance after 50k InterSyn fine-tuning. Values in parentheses denote the change ($\Delta$) from the base.
Model MME-P MMBench MMMU SEEDBench
VILA-U 1344 (+8) 57.1 (+0.8)
VARGPT 1465 (-23) 66.8 (-0.8) 37.2 (+0.8) 65.6 (-2.3)
VARGPT-v1.1 1658 (-26) 79.4 (-1.6) 46.2 (-2.3) 75.2 (-0.9)
BAGEL 1646 (-41) 83.1 (-1.9) 52.8 (-2.5)
Crucially, the substantial gains in interleaved generation capabilities (demonstrated in §4.2.1) do not come at the cost of core understanding performance. As shown in Table 2, the models’ performance on standard understanding benchmarks (Fu et al., 2023; Liu et al., 2025; Yue et al., 2024; Li et al., 2024) remains robust after fine-tuning.
4.2.3 Effectiveness of Multi-turn Data
To verify the effectiveness of our multi-turn dialogue data, we designed an experiment to assess its impact on a models’ conversational capabilities. We fine-tuned two models capable of multi-turn generation, Anole and VARGPT-v1.1, on different compositions of single-turn and multi-turn data, while keeping the total training size fixed at 50k samples. The goal is to demonstrate that training with multi-turn data enhances a model’s ability to maintain context and quality across an extended conversation. The performance, evaluated by TCC, ICC, IQ, and ITS, is presented in Table 3.
Table 3: Effectiveness of multi-turn data on conversational performance across across different dialogue test turns. Models are trained on different proportions of single-turn and multi-turn data.
Model Setting Training Data Turn 1 Turn 2 Turn 3
Single Multi TCC ICC IQ ITS TCC ICC IQ ITS TCC ICC IQ ITS
Anole Baseline - - 3.09 3.01 2.92 2.26 2.80 2.75 2.60 1.90 2.40 2.30 2.10 1.40
Trained 50k 0 3.47 3.28 3.10 2.74 3.25 2.85 2.70 2.40 2.85 2.60 2.45 2.05
Trained 25k 25k 3.48 3.27 3.13 2.77 3.35 3.00 2.80 2.40 3.00 2.70 2.55 2.10
Trained 0 50k 3.52 3.24 3.10 2.94 3.40 3.05 2.90 2.55 3.27 2.85 2.70 2.25
VARGPT-v1.1 Baseline - - 3.26 1.01 1.23 0.68 3.10 0.95 1.18 0.72 2.90 0.97 0.90 0.65
Trained 50k 0 3.68 3.12 3.67 3.00 3.40 2.90 3.45 2.90 3.05 2.60 3.05 2.65
Trained 25k 25k 3.65 3.18 3.66 3.18 3.45 2.95 3.40 2.90 3.21 2.78 3.24 2.66
Trained 0 50k 3.64 3.20 3.68 3.11 3.58 3.10 3.52 3.05 3.48 2.95 3.45 2.90
The results clearly demonstrate the value of multi-turn training data. First, the inclusion of multi-turn data does not compromise first-turn performance, which remains high across all trained settings. This is expected, as the SEIR pipeline ensures comparable data quality regardless of dialogue type. Second, and more importantly, multi-turn training significantly reduces performance degradation in later turns. Models trained only on single-turn data show a steeper drop in quality as the conversation progresses. In contrast, training with multi-turn dialogues mitigates this degradation, particularly for ITS, by explicitly teaching the model to maintain context across extended conversational dependencies. This confirms the effectiveness of our multi-turn dataset in fostering more coherent and consistent multi-turn generation.
4.3 Effectiveness of SEIR
To validate the effectiveness of the SEIR, we conduct experiments on both the iterative refinement process and the final output quality comparison with other generators.
4.3.1 Validation of Iteration Refinement
In this section, all evaluations in this section are conducted by human judge to establish a ground truth. For question quality, Figure 3 shows that QR improves quality over the first three iterations but plateaus thereafter. Based on this, we set the QR to 3, achieving 99.5% of peak quality while reducing computational cost by 40%. For answer quality, we evaluate across different iterations of AR and IR, using a set of 7,000 questions generated through three rounds of QR. As shown in Table 4.3.1, the results confirm that AR primarily improves content completeness, while IR enhances multimodal synergy, demonstrating the effectiveness of the SEIR method. Experimental results show that when both AR and IR are set to 4 or 5, the improvements become marginal. Based on this, we fix the number of AR and IR iterations to 3 in dataset construction settings.
@@ 4.3.1 Validation of Iteration Refinement
Table 4: Impact of answer refinement (AR) and image refinement (IR) on answer quality. The table reports human evaluation mean scores across four dimensions (TCC, ICC, IQ, ITS). AR improves TCC, ICC, and ITS, while IR further enhances ICC and ITS, confirming the effectiveness of iterative refinement.
@@
4.3.2 Comparison of SEIR Outputs with Other Generators
We evaluate the SEIR pipeline against 13 baseline generators using the fixed Evaluation Benchmark defined in §3.5.3. In this section, we employ both human judges and SynJudge to assess the generator outputs across the four evaluation dimensions.
@@ 4.3.2 Comparison of SEIR Outputs with Other Generators
As shown in Table 5, both human judges and SynJudge confirm that SEIR-generated samples (InterSyn) achieve the highest mean scores across all dimensions. They outperform the strongest baseline, GPT-4o+DALL-E, by a margin of 0.34–0.66, with the largest gap observed in ITS. Furthermore, InterSyn exhibits very low variance (below 0.61), reflecting its consistent output quality and the robustness of our automated method. Crucially, the significant performance gap between InterSyn and the best SOTA generators reveals that even top models still struggle with image–text alignment and complementarity, indicating substantial room for future improvement.
Table 5: Generator performance evaluated by human judge and SynJudge. Each entry is reported as mean (variance), where the value outside the parentheses denotes the mean score and the value inside the parentheses denotes the variance.
Generator Human SynJudge
TCC ICC IQ ITS TCC ICC IQ ITS
Anole 3.06 (1.47) 2.95 (2.05) 2.89 (1.87) 2.25 (2.47) 3.09 (1.47) 3.01 (2.11) 2.92 (1.93) 2.26 (2.55)
DDiT 0.38 (1.17) 3.51 (0.83) 3.29 (0.85) 0.37 (1.21) 0.28 (0.86) 3.67 (0.75) 3.34 (0.92) 0.26 (0.87)
Emu3 3.38 (0.86) 3.86 (0.86) 3.87 (0.59) 3.26 (1.31) 3.37 (0.74) 3.92 (0.66) 3.85 (0.80) 3.37 (1.16)
VILA-U 2.45 (2.58) 3.62 (0.91) 3.35 (0.97) 2.31 (3.01) 2.46 (2.73) 3.72 (0.92) 3.37 (1.12) 2.19 (2.96)
Liquid 2.88 (0.76) 3.82 (0.69) 3.67 (0.70) 3.00 (1.42) 2.87 (0.71) 3.86 (0.75) 3.76 (0.85) 3.06 (1.36)
Janus-Pro 2.93 (0.95) 3.25 (1.21) 3.16 (0.99) 2.55 (1.75) 2.96 (0.96) 3.30 (1.09) 3.11 (1.11) 2.62 (1.87)
Show-o 3.32 (1.31) 3.65 (0.94) 3.52 (0.89) 3.10 (1.94) 3.49 (1.12) 3.79 (0.82) 3.57 (0.96) 3.30 (1.70)
Show-o-Turbo 3.48 (1.12) 3.77 (0.88) 3.53 (0.94) 3.33 (1.60) 3.50 (1.06) 3.89 (0.86) 3.64 (1.01) 3.45 (1.44)
VARGPT 2.60 (0.66) 0.94 (2.37) 0.94 (2.32) 0.55 (1.99) 2.55 (0.30) 0.94 (2.39) 0.87 (2.17) 0.67 (2.16)
VARGPT-v1.1 3.13(0.83) 0.89(1.98) 1.16(2.11) 0.72(1.83) 3.26(0.76) 1.01(2.12) 1.23(1.95) 0.68(1.96)
BAGEL 2.97(0.83) 3.92(0.81) 4.18(0.75) 2.81(1.21) 3.11(0.91) 3.89(0.72) 4.23(0.66) 2.87(1.33)
Gemini+Flux 3.94 (0.94) 4.06 (0.58) 4.43 (0.47) 3.81 (0.90) 3.97 (0.57) 4.12 (0.71) 4.48 (0.64) 3.84 (1.11)
GPT-4o+DALL-E 4.05 (0.37) 4.08 (0.48) 4.41 (0.57) 3.94 (0.64) 3.99 (0.65) 4.10 (0.81) 4.45 (0.58) 3.87 (1.16)
SEIR 4.41 (0.55) 4.46 (0.55) 4.47 (0.53) 4.51 (0.57) 4.39 (0.61) 4.49 (0.63) 4.44 (0.45) 4.53 (0.51)
4.4 Reliability of SynJudge
4.4.1 Evaluation Setup for Judges
To identify a reliable automatic evaluator that aligns closely with human scoring, we conduct a comparative experiment. The protocol is designed to rigorously measure each candidate judge’s deviation from a human-annotated ground truth.
We evaluate a total of five model-based judges against our human ground truth. The candidates are: (1) Human Judge (Ground Truth): A panel of ten experts whose scores serve as the gold standard. A cross-review protocol was used to ensure scoring reliability and mitigate individual bias. (2) Zero-Shot MLLM Judges: Three off-the-shelf MLLMs used for automated assessment: GPT-4o (OpenAI, 2024), QwenVL2.5 (Bai et al., 2025), and InternVL2.5 (Chen et al., 2024c). (3) SynJudge Candidates (Finetuned): We fine-tuned two strong MLLM backbone candidates, QwenVL-trained and InternVL-trained, to create our proposed evaluator.
The evaluation is conducted on a test set of 9,600 human-annotated question-answer pairs. To provide a comprehensive assessment of judge performance, we use two complementary metrics: (1) Root Mean Squared Error (RMSE): Measures the magnitude of the deviation from human scores. (2) Human Agreement (A@1): Measures the percentage of scores that are within a 1-point tolerance of the human rating, reflecting practical reliability. A lower RMSE and a higher A@1 indicate stronger alignment with human judgment. The SynJudge candidates were trained on a separate training set of 38,400 human-annotated pairs. Detailed definitions of the metrics and training hyperparameters of SynJudge are provided in the appendix (§D.2, §D.4).
4.4.2 Comparison Results
Figure 4 and Table 6 report the performance of each judge using both RMSE and A@1 metrics. The results show a clear trend: finetuned judges decisively outperform zero-shot models on both metrics.
The QwenVL-trained judge demonstrates the strongest alignment with human preferences, achieving both the lowest average RMSE and the highest average A@1 of 95.4%. This high agreement rate signifies that over 95% of its scores are within one point of human judgment, confirming its high reliability. In contrast, zero-shot judges like GPT-4o lag significantly, with A@1 scores around 86.5%. Based on this superior performance across complementary metrics, we select QwenVL-trained as our final SynJudge.
Figure 4: Visualization of RMSE for different judges.
Table 6: Judge performance comparison. We report average RMSE (lower is better) and Human Agreement (A@1, higher is better) against human scores. The best result in each row is highlighted in bold. QwenVL_trained demonstrates the strongest alignment.
Dimension GPT-4o QwenVL InternVL QwenVL_trained InternVL_trained
TCC (RMSE) 1.01 0.81 0.96 0.54 0.55
ICC (RMSE) 1.02 1.09 0.90 0.72 0.70
IQ (RMSE) 0.98 0.96 1.06 0.68 0.72
ITS (RMSE) 1.18 1.20 1.03 0.67 0.72
A@1 0.865 0.875 0.866 0.954 0.945